OpenStudy (anonymous):

According to one study, 61% of the population swallow at least one spider per year in their sleep. Based on this study, what is the probability that exactly 7 of 10 randomly selected people have swallowed at least one spider in their sleep in the last year?

OpenStudy (dumbcow):

use binomial distribution \[P(X=k) = \left(\begin{matrix}n \\ k\end{matrix}\right)p^{k}(1-p)^{n-k}\] n=10 k = 7 p = 0.61